OpenStudy (anonymous):

At elevated temperatures, dinitrogen pentoxide decomposes to nitrogen dioxide and oxygen: 2N 2 O 5 (g) 4NO 2 (g) + O 2 (g) When the rate of formation of O 2 is 2.2 × 10 - 4 M/s, the rate of decomposition of N 2 O 5 is __________ M/s.

OpenStudy (anonymous):

please I need help with this worksheet

OpenStudy (cuanchi):

for each mole of O2 that is form in the reaction 2 moles of N2O5 has to be consumed then if you have the reaction produced 2.2 x 10^-4 M/s of O2 you wil be able to figure out how much N2O5 has to be consumed by second to produce that amount of O2.

OpenStudy (cuanchi):

\[Rate= -\frac{ \Delta[N _{2}O _{5}] }{2 \Delta t}=\frac{ \Delta[O _{2}] }{ \Delta t}\] \[-\frac{ \Delta[N _{2}O _{5}] }{\Delta t}=\frac{ 2\Delta[O _{2}] }{ \Delta t}\] \[-{ \Delta[N _{2}O _{5}] }={ 2\Delta[O _{2}] }\]